A voltaic cell uses the following reaction: 2Al(s) + 3I2(s) → 2Al3+(aq) + 6I-(aq)

a) Calculate the cell potential (ε°cell) under standard conditions.

b) Calculate the cell potential (E) when [Al3+] = 0.015M and [I-] = 0.025M.
Answer #1

a) E tu a to ako - El Ox = Efele - El portal de = 0.536-61-662) = (2-1884 b) NERNES EQn. Freue = txu - O.ospe long | Latesty?